OpenStudy (anonymous):

Solve the following system of equations. Show Me How To Do This Please..! ^_^ 4x + 3y = -1 3x + y = 3

OpenStudy (anonymous):

What methods have you learned so far?

OpenStudy (anonymous):

The method of elimination and substitution.

OpenStudy (anonymous):

u can use them, elimination is awesome.

OpenStudy (anonymous):

Let's try substitution. From here I can tell that the easiest thing we can so is substitute the y from the second equation into the first one but we want to isolate y first.

OpenStudy (anonymous):

\[3x + y = 3 \]\[3x-3x + y = 3-3x \]\[ y = 3-3x \]

OpenStudy (anonymous):

Now we substitute the equation to the first one which is \[4x + 3y = -1\]

OpenStudy (anonymous):

okay yes...Then what..?

OpenStudy (anonymous):

That then becomes \[4x + 3(3-3x) = -1\] the y disappears

OpenStudy (anonymous):

now solve that equation for x

OpenStudy (anonymous):

Can you tell me what you get?

OpenStudy (anonymous):

first lets eliminate x: multiply both eqtns by a number that will make the value of their Xs equal: (4x+3y= -1) 3 (3x + y= 3) 4 we will have new sets of eqtns: 12x+9y=-3 (-) 12x +4y=12 subtracting, wecan eliminate x and the answer will be: 5y= -15 therefore, y = -3 substituting the value of y to any of the eqtn, we can get: 3x + (-3) = 3 x = 1 to check, just substitute the values of x and y. :)

OpenStudy (anonymous):

okay thank you